### v1

@@ 3.5.2. Attractive and Repulsive Weighted HK models
If we run a simulation with 100 normal agents, a manipulative group with $K\in\{5j\}_{j=0}^{30}$ agents who change their opinions from $x=-0.8$ to $x=0.8$ in $t_{\Delta}\in\{5i\}_{j=1}^{30}$ iterations, we obtain the graph in figures 22 and 23.
Figure 22. Simultaneous Repulsive Weighted HK model. After 100 simulations with $\varepsilon=0.1$ and 100 normal agents with a manipulative group with $K\in\{5j\}_{j=0}^{30}$ agents changing their opinions from $f(0)=-0.8$ to $f(t_{\Delta})=0.8$ in $t_{\Delta}\in\{5j\}_{j=1}^{30}$ iterations. The graphs show the averages of the mean opinions (left) and the average of final opinions (right).
Figure 23. Simultaneous Repulsive Weighted HK model. Simulation with the same parameters as in figure 22 except the confidence threshold, which is constant $\varepsilon=0.6$.
Observe that, for $\varepsilon=0.1$ and $t=t_{\Delta}$ iterations, the effect of the manipulative group’s action increases approximately to $\overline{\mathbf{x}}=0.35$ the average of the normal opinions when the values of iterations $t_{\Delta}$ and the number of manipulators $K$ are large (yellows in figure 22, top left). Recall that, for this simulation, we have chosen values of $K$ up to 1.5 times the number of normal population $N=100$. Bearing in mind that, without the influence of the manipulators, the opinions are grouped into two clusters of (approximately) the same size at $x=\pm 1$. Let us call $\mu$ the number of normal agents that pass from the group that converges to $x=-1$ to the group that converges to $x=1$. When a manipulative group acts obtaining the same two final clusters at $x=\pm 1$, we can compute the average
@@
then if the average satisfies $0\leq\overline{\mathbf{x}}\leq 0.35$, then $0\leq\mu\leq 17$. This implies that a manipulative group with 150 agents is not able to modify the opinion of more than 17 normal agents (17 % of the total normal population).
The dispersion of opinions at $t=t_{\Delta}$ decreases as $t_{\Delta}$ increases, because when this value is small, the number of iterations is not large enough for the dynamics to stabilize (figure 22, top right). This stabilization can be reached in the range $t\in[t_{\Delta},1000]$ when the manipulative group behaves as stubborn.
If we compare with the results at $t=1000$, the final distributions consist of two big groups at $x=\pm 1$ and some isolated agents with moderated opinions. In addition, as in the attractive model, for small values of $t_{\Delta}$, the value of the mean moves in the opposite direction to the movement of the manipulator group (in dark blue in figure 22, bottom left) with a major effect reaching $x=-0.35$.
In the case of $\varepsilon=0.6$ (figure 23), we observe that the average of normal opinions at $t=t_{\Delta}$ is, approximately, in the range $0.20\leq\overline{\mathbf{x}}<$ 0.2 and the distribution of the normal opinion consists, mainly, in two clusters at $x=\pm 1$ for values $t_{\Delta}\geq 50$ (dark blue in figure 23, top right). When we consider $t=1000$ iterations, the final distribution is bimodal $x=\pm 1$ for any values of $t_{\Delta}$ and $K$.
3.5.3. Attractive–Repulsive Weighted HK model
@@
We have seen in section 2.3 that the dynamics of the simultaneous Attractive–Repulsive Weighted HK model, for small values of $\varepsilon$, consists of a set of opinions oscillating in the central range of the opinion space, while extreme opinions $x=\pm 1$ can also group close normal agents together. Some central oscillating opinions can escape to converge to one of those extremal clusters.
@@
With the intervention of manipulators, normal opinions outside the confidence interval of the manipulative group are moved away from this group. Because this interval is small, the opinions that the manipulative group can attract are few, and this attraction can be maintained throughout the iterations when the variation of the manipulative opinions is slow, that is, when $t_{\Delta}$ is large.
@@
Thus, in the simulation with 100 normal agents, a manipulative group of 80 agents whose opinions vary from $f(0)=-0.8$ to $f(t_{\Delta})=0.8$ and a confidence threshold of $\varepsilon=0.1$, it is observed (cf. figure 24) that, as the iterations run to $t=t_{\Delta}$, the manipulative group presses the normal opinions with higher values towards $x=1$, making their convergence to this cluster faster. On the other hand, the range of opinions lower than that of the manipulators becomes wider and allows normal opinions to oscillate. This oscillation will be displaced downwards by the action of the repulsive mechanism when the manipulative group becomes stubborn. As $t_{\Delta}$ gets bigger, these oscillating opinions eventually converge to cluster $x=-1$ (cf. figure 24, center and right).
@@
Figure 24. Simultaneous Attractive–Repulsive Weighted HK model. Simulation with 100 normal agents interacting with a group of 80 manipulators with opinions initial $f(0)=-0.8$ and final $f(t_{\Delta})=0.8$ reached in $t_{\Delta}\in\{20,40,60\}$ iterations. The confidence threshold is fixed at $\varepsilon=0.1$ constant.
@@
To study the dependence of the mean opinions on the number of manipulators $K$ and the number of iterations $t_{\Delta}$ during which the manipulative group varies its opinion, we will consider simulations for $\varepsilon\in\{0.1,0.6\}$, analogous to those in the previous section. Thus, we will set $N=100$ the number of normal agents, while the variables $t_{\Delta}$ and $K$ take values in the ranges $t_{\Delta}\in\{5j\}_{j=1}^{40}$ and $K\in\{5j\}_{j=0}^{40}$. In this case, we have chosen to extend the ranges of $t_{\Delta}$ and $K$ in order to better observe the changes in mean opinions.
@@
For the confidence threshold $\varepsilon=0.1$, we can notice that the average of the mean opinions at iteration $t=t_{\Delta}$ (see figure 26 top left) reaches approximately the value $x\simeq 0.5$ if $K\simeq 200$ and $t_{\Delta}\simeq 200$, but remains close to $x=0$ when $K\lesssim 80$ (in blue colors). The mean opinion will increase as the value of $K$ increases, and specially from $K\simeq 80$, and favouring this increase in the mean opinion, the large values of $t_{\Delta}$ over the small ones.
@@
It is important to note that this fact implies that, in order for the manipulative group to have a significant influence on the average opinion, the size of the group must be greater than 80% of the size of the normal population, unlike what happens in the HK and DW models, in which a smaller manipulative group can have a much more important influence.
@@
Again, because of the tendency to polarization in a bimodal distribution at $x=\pm 1$, even with the existence of moderate oscillating opinions, the average of the standard deviations is large and the number of clusters approaches 2 (in dark blue in figure 26 at the top right) when the values of $t_{\Delta}$ cease to be small.
@@
Figure 26. Simultaneous Attractive–Repulsive Weighted HK model. After 100 simulations with $\varepsilon=0.1$ and 100 normal agents with a manipulative group with $K\in\{5j\}_{j=0}^{40}$ agents changing their opinions from $f(0)=-0.8$ to $f(t_{\Delta})=0.8$ in $t_{\Delta}\in\{5j\}_{j=1}^{40}$ iterations. The graphs show the averages of the mean opinions (left), the standard deviation (centre) and the number of final opinions (right).
@@
In the bottom row of figure 26, we show the results after $t=500$ iterations in such a way that the opinion of the manipulative group is $f(t)=0.8$ for all $t\in\{t_{\Delta},\ldots,500\}$, that is, the manipulators become stubborn agents. In this case, oscillating opinions (if any) tend to converge to one of the extreme clusters (or may remain oscillating), stabilizing the average values studied.
@@
Thus, when $t_{\Delta}$ is small, the averages of the mean opinions move in the opposite direction to the opinion of the manipulative group, even reaching $x\simeq-0.25$. Again, this is because the manipulative group changes its opinion quickly and before the normal opinions converge to one of the clusters, and since the confidence interval is small, the repulsive mechanism pushes the normal opinions in the opposite direction to that of the manipulators (see dark blue region in the figure 26 bottom left). The stabilization of the values is also observed by the reduction in the final number of clusters, which are between 2 and 9 (cf. figure 26 bottom right).
@@
When we set the confidence threshold at $\varepsilon=0.6$, the attractive mechanism has greater influence and the manipulative group is able to attract a greater number of normal opinions. After $t=t_{\Delta}$ iterations (see top row of figure 27), we can observe that the highest average of the mean opinions is reached for values $(t_{\Delta},K)\simeq(40,200)$ (in yellow in figure 27 top left). For greater values of $t_{\Delta}$, the manipulative group needs more iterations to reach its final opinion. Thus, normal opinions in the confidence interval of the manipulative group are more likely to be repelled by other agents while they follow the manipulative group (seen in figure 25).
It is also observed, even more clearly than in the case of $\varepsilon=0.1$, that when $K\lesssim 60$, the influence of the manipulating group is small (note clearly delimited the dark blue region in figure 27 above left). The behaviour of the average of the final number of clusters is similar to the $\varepsilon=0.1$ case, in which, only for small values of $t_{\Delta}$, there are a large number of opinions that are not yet clustered.
When we consider the dynamics after $t=500$ iterations, we observe that the average of the mean opinions remains similar except for small values of $t_{\Delta}$, for which the effect of the shift in the opposite direction to that of the manipulators’ opinion is repeated. As for the number of clusters, the graph at the bottom right of figure 27 shows, in dark blue, a region in which the final distribution converges to the extreme values $x=\pm 1$, except in the yellow region, which corresponds to a distribution with a large number of opinions that oscillate attracted by the cluster $x=1$ and the manipulative group (now stubborn) at $x=0.8$ as seen in figure 25 left. This region of oscillating opinions is clearly delimited as figure shows.
Figure 27. Simultaneous Attractive–Repulsive Weighted HK model. Simulation with the same parameters as in figure 26 except the confidence threshold, which is constant $\varepsilon=0.6$.
@@ 4. Conclusions and discussion
We have studied the opinion dynamics in various models of bounded confidence under the influence of one group of manipulative agents. This type of agent, as a generalization or extension of stubborn agents, allows them to have influence within the union of the confidence intervals corresponding to their range of opinions. Thus, strategies can be established to modify the natural dynamics of the model (due only to normal agents). This modification can cause opinions acceptable to society of normal agents to become extreme or vice versa. The range of opinions can be moved as well as expand or contract in the space of opinions, thus reminiscent of the Overton window theory. Therefore, manipulators may attempt to modify the Overton window to form a new framework of acceptable opinions.
@@
In this article, we have described the strategies of one single manipulative group that varies its opinions linearly over a wide range of opinions, and then becomes a stubborn group to the point of convergence or stabilization of the dynamics. The number of manipulators in the group, as well as the rate of variation in their opinions, have an unequal influence on the different models studied, with the different confidence thresholds. Due to the arbitrariness of choosing a canonical definition for the Overton window, we consider that the mean as the central point and the standard deviation as the amplitude can be used to determine it. We have shown examples of simulations of opinion dynamics that allow us to interpret the results of the change in the Overton window, using mean values and standard deviations, based on the variables that define the manipulative group. These changes in values resulting from dynamics are quantified with a group of manipulators that move in a wide and fixed range for the different models. In the HK and DW models, because of their tendency to central clustering, a manipulative group with a sufficient number of agents and/or whose change of opinion is sufficiently slow, can attract a large number of normal agents and take them with them to any point in the opinion space. Moreover, for the simultaneous HK model, the conditions of these variables have been fully determined so that manipulators attract the opinions of another group of normal agents. In Weighted HK models, a larger number of manipulators is needed to influence the opinion of normal agents, and rapid changes in the opinion of manipulators could have the opposite effect on normal agents.
@@
There are still pending issues that will be studied in future articles, among them, the competition between various groups of manipulators according to their size and the trajectories of their opinions, the finding of new models that adapt to the movement of the Overton window placing it in the central point of the space of opinions. In this sense, the density function $\phi_{i}(t)=1-|x_{i}(t)|$ in the iteration of the Weighted HK model can be modified so that its mean (or median) in each iteration matches the mean (or median) of the distribution of normal opinions.
@@ Acknowledgements
The author would like to thank Gemma Rioja for her invaluable suggestions and comments that have substantially improved the text of this article.
@@
All the simulations and graphs in this article have been made using the R2020b version of Matlab.

### v2

@@ 5.1.3. Repulsive Weighted HK model with manipulators
In figure 21, we can observe that for $\varepsilon=0.1$ and $t_{\Delta}>0$, the effect of the manipulative group increases the mean values in the range of interval $[0,0.55]$ at iteration $t=t_{\Delta}$. On the other hand, at $t=1000$, when the iteration converges to the bimodal distribution $x=\pm 1$, the mean values are in $[-0.2,0.55]$ obtaining the negative values for small values of $t_{\Delta}$ and $K$. This shows that if the manipulative group changes its opinion quickly, it will get the opposite result to the one intended.
Figure 21. Simultaneous Repulsive Weighted HK model: averages of mean opinions (left), of standard deviations (center) and of number of final clusters (right) after 100 simulations for every $K\in\{5j\}_{j=0}^{40}$ and $t_{\Delta}\in\{5j\}_{j=1}^{40}$. The manipulative group varies its opinion from $f(0)=-0.9$ to $f(t_{\Delta})=0.9$. The confidence threshold is fixed at $\varepsilon=0.1$ constant.
Since the final distribution converges to $x=\pm 1$, we can call $\mu$ the difference between the number of normal agents converging to $x=1$ and those converging to $x=-1$. Then we can compute the average
@@
then if the average satisfies $0\leq\overline{\mathbf{x}}\leq 0.55$, then $0\leq\mu\leq 28$. This implies that a manipulative group with 200 agents is not able to modify the opinion of more than 28 normal agents (28 % of the total normal population).
In the case of $\varepsilon=0.6$, the dynamics show similar behaviour but with less ability of the manipulators to modify the opinion of the normal agents.
Notice that, since the final distribution converges to the values $x=\pm 1$ then they are primary clusters and the interval defined by those is $[-1,1]$. So, this metric do not offer further information for this model.
5.1.4. Attractive–Repulsive Weighted HK model with manipulators
@@
In section 5, we have seen that the dynamics of the simultaneous Attractive–Repulsive Weighted HK model, for small values of $\varepsilon$, consists of a set of opinions oscillating in the central range of the opinion space, while extreme opinions $x=\pm 1$ can also group close normal agents together (see figure 15). Some central oscillating opinions can escape to converge to one of those extremal clusters.
@@
With the intervention of manipulators, normal opinions outside the confidence interval of the manipulative group are moved away from this group. When this interval is small, the normal opinions that the manipulative group can attract are few, and this attraction can be maintained throughout the iterations when the variation of the manipulative opinions is slow, that is, when $t_{\Delta}$ is large.
@@
Thus, in the simulation with 100 normal agents, a manipulative group of 80 agents whose opinions vary from $f(0)=-0.9$ to $f(t_{\Delta})=0.9$ and a confidence threshold of $\varepsilon=0.1$, it is observed (cf. figure 22) that, as the iterations run to $t=t_{\Delta}$, the manipulative group presses the normal opinions with higher values towards $x=1$, making their convergence to this cluster faster. On the other hand, the range of opinions lower than that of the manipulators becomes wider and allows normal opinions to oscillate. This oscillation will be displaced downwards by the action of the repulsive mechanism when the manipulative group becomes stubborn. As $t_{\Delta}$ gets bigger, these oscillating opinions eventually converge to cluster $x=-1$ (cf. figure 22, center and right).
@@
Figure 22. Simultaneous Attractive–Repulsive Weighted HK model: simulation with 100 normal agents interacting with a group of 80 manipulators varying its opinion from $f(0)=-0.9$ to $f(t_{\Delta})=0.9$ for $t_{\Delta}\in\{20,40,60\}$ iterations (dashed line). The confidence threshold is fixed at $\varepsilon=0.1$ constant.
@@
5.1.5. Simulations for Attractive–Repulsive Weighted HK model with one manipulative group
Now, we will consider $N=100$ equispaced normal agents in $[-1,1]$, while the variables $t_{\Delta}$ and $K$ take values in the ranges $t_{\Delta}\in\{5j\}_{j=0}^{40}$ and $K\in\{5j\}_{j=0}^{40}$.
@@ 5.1.5. Simulations for Attractive–Repulsive Weighted HK model with one manipulative group
For the confidence threshold $\varepsilon=0.1$ (see figure 24), we can notice that the average of the mean opinions measured at iteration $t=t_{\Delta}$ reaches approximately the value $x\simeq 0.5$ for large values of $K$ and $t_{\Delta}$, but rarely reaches $x=0.2$ when $K\lesssim 75$.
@@
It is important to note that the manipulative group has a significant influence on the average opinion when the size of the group is greater than 75% of the size of the normal population, contrarily what happens in the HK and DW models, in which smaller manipulative groups can have a much more important influence.
@@
Because the tendency to polarization in a bimodal distribution at $x=\pm 1$, even with the existence of moderate oscillating opinions, the average of the standard deviations is large and the number of clusters approaches 2 when the values of $t_{\Delta}$ cease to be small.
@@
Figure 24. Simultaneous Attractive–Repulsive Weighted HK model: averages of mean opinions (left), of standard deviations (center) and of number of final clusters (right) after 100 simulations for every $K\in\{5j\}_{j=0}^{40}$ and $t_{\Delta}\in\{5j\}_{j=1}^{40}$. The manipulative group varies its opinion from $f(0)=-0.9$ to $f(t_{\Delta})=0.9$. The confidence threshold is fixed at $\varepsilon=0.1$ constant.
@@
In the bottom row of figure 24, we show the results after $t=500$ iterations in such a way that the manipulative group becomes stubborn at $f(t)=0.9$ for all $t\in\{t_{\Delta},\ldots,500\}$. In this case, oscillating opinions (if any) tend to converge to one of the extreme clusters (or may remain oscillating), stabilizing the average values studied.
@@
Observe that when $t_{\Delta}\leq 50$ and $K\gtrsim 100$, the averages of the mean opinions move in the opposite direction to the opinion of the manipulative group, even reaching $x\simeq-0.7$ (see dark blue region in the figure 24 bottom left). Again, this is because the manipulative group changes its opinion quickly and before the normal opinions converge to one of the clusters, and since the confidence interval is small, the repulsive mechanism pushes the normal opinions in the opposite direction to that of the manipulators. The stabilization of the values is also observed by the reduction in the final number of clusters at $t=500$ in relation with $t=t_{\Delta}$ (see figure 26).
Figure 25. Simultaneous Attractive–Repulsive Weighted HK model: histograms of the effective weights in the simulation of figure 24. The most frequent effective weights are those closest to zero.
There is not an adequate cluster screening by their effective weights because the most frequent are the effective weights closest to $W\simeq 0$ (see figure 25). If we consider that each cluster is primary (filtering with the screening value $\delta=0$) then we obtain that the range of effective weights is almost the entire interval $[-1,1]$. When we choose $\delta=0.2$ for screening, we can have only one primary cluster losing the extreme cluster $x=1$ close to the opinion of the manipulators.
Figure 26. Simultaneous Attractive–Repulsive Weighted HK model: averages of centers, amplitudes and number of number of clusters in the interval of primary clusters with $\delta=0$ of the same simulation of figure 24.
@@
When we set the confidence threshold at $\varepsilon=0.6$, the attractive mechanism has greater influence and the manipulative group is able to attract a greater number of normal opinions. Notice that, after $t=t_{\Delta}$ iterations, the highest average of the mean opinions is reached for values $(t_{\Delta},K)\simeq(50,200)$ (in yellow in figure 27 top left). For greater values of $t_{\Delta}$, the manipulative group needs more iterations to reach its final opinion. Thus, normal opinions in the confidence interval of the manipulative group are more likely to be repelled by other agents while they follow the manipulative group.
It is also observed, that the influence of the manipulating group is smaller that in the case $\epsilon=0.1$. In fact, the rejection effect that we found for the case $\epsilon=0.1$ does not occur, in which the opinions shifted in the opposite direction to that of the manipulating group.
Figure 27. Simultaneous Attractive–Repulsive Weighted HK model: averages of mean opinions (left), of standard deviations (center) and of number of final clusters (right) after 100 simulations for every $K\in\{5j\}_{j=0}^{40}$ and $t_{\Delta}\in\{5j\}_{j=1}^{40}$. The manipulative group varies its opinion from $f(0)=-0.9$ to $f(t_{\Delta})=0.9$. The confidence threshold is fixed at $\varepsilon=0.6$ constant.
Now, it is possible to screen by primary clusters since the most frequent effective weights are close to $W=1.2$ (see histogram in figure 28).
Figure 28. Simultaneous Attractive–Repulsive Weighted HK model: histograms of the effective weights in the simulation of figure 27.
If we choose the screening value $\delta=0.5$ at the iteration $t=t_{\Delta}$, we observe that for $K\leq 100$, we get two primary clusters at $x=\pm 1$ and only for values of $K>150$ and $t_{\Delta}$ small, we obtain an only primary cluster (yellow region at top left of figure 29). But, in this region, the final distribution at $t=500$ iterations converges to two primary clusters.
Figure 29. Simultaneous Attractive–Repulsive Weighted HK model: average of centers, amplitudes and number of number of clusters in the interval of primary clusters with $\delta=0.5$ of the same simulation of figure 27.
@@ 6. Conclusions and discussion
We have studied the opinion dynamics in various models of bounded confidence under the influence of one group of manipulative agents. This type of agent, as a generalization or extension of stubborn agents, allows them to have influence within the union of the confidence intervals corresponding to their range of opinions. Thus, strategies can be set up to modify the natural dynamics of the model (due only to normal agents). This modification can cause socially acceptable opinions of normal agents to become extreme or vice versa. The range of opinions can shift as well as expand or contract in the opinion space, reminiscent of Overton window theory. Therefore, manipulators can try to modify the Overton window to form a new framework of acceptable opinions.
@@
In this article, we have described the strategies of one single manipulative group that linearly varies its opinions over a wide range of opinions, and then becomes a stubborn group until convergence or stabilization of the dynamics. The number of manipulators in the group, as well as the rate of variation of their opinions, have an unequal influence on the different models studied, with the different confidence thresholds. Due to the arbitrariness of choosing a definition for the Overton window, we consider that the mean as the central point and the standard deviation as the width can be used to determine it. In some cases, when possible, we have used the center and amplitude of the interval of primary clusters to measure the influence of manipulators, with similar results. We have shown examples of simulations of opinion dynamics with a group of manipulators interpreting and quantifying their influence on normal agents as a function of the variables defining that group. In the HK and DW models, due to their tendency to central clustering, a manipulative group with a sufficient number of agents and/or whose change of opinion is sufficiently slow, can attract a large number of normal agents and take them with them anywhere in the opinion space. Moreover, for the simultaneous HK–model, the conditions of these variables have been fully determined so that manipulators attract the opinions of another group of normal agents. In Weighted HK models, a larger number of manipulators is needed to influence the opinion of normal agents, and rapid changes in the opinion of manipulators could have the opposite effect on normal agents.
@@
There are still pending issues that will be studied in future articles, among them, the competition between various groups of manipulators according to their size and the trajectories of their opinions, the finding of new models that adapt to the movement of the Overton window placing it in the central point of the space of opinions. In this sense, the density function $\phi_{i}(t)=1-|x_{i}(t)|$ in the iteration of the Weighted HK model can be modified so that its mean in each iteration matches the central point (mean or center) of the distribution of normal opinions.
All the simulations and graphs in this article have been made using the R2020b and R2024a versions of Matlab.
@@ Acknowledgements
The author would like to thank the two anonymous referees for their comments, especially on the structuring of the article and for the incorporation of the primary clusters metric, and also Gemma Rioja for the suggestions provided on the wording. These contributions have significantly improved this article.